Many single-celled organisms propel themselves through the water with long tails, which they wiggle back and forth. (The most obvious example is the sperm cell.) The frequency of the tail's vibration is typically about 10-15 Hz. To what range of periods does this range of frequencies correspond?
Answer #1

fL = lower frequency level = 10 Hz

fU = Upper frequency level = 15 Hz

we know that , time period = 1/frequency

So TH = 1/fL = 1/10 = 0.1 sec

and

TU = 1/fU = 1/15 = 0.067 s

So range of period is 0.067 - 0.1